Please help me with this problem:
(2t-5)^2=24
I think the finished problem is written with a sqrt.

Answer by stanbon(75887)

Please help me with this problem:
(2t-5)^2=24
I think the finished problem is written with a sqrt.
---------------
2t-5 = sqrt(24) or 2t-r = -sqrt(24)
----
t = (2sqrt(6)+5)/2 or t = (-2sqrt(6)+5)/2
=============
Cheers,
Stan H.
=============